OpenStudy (anonymous):

Please answer this>>> i get e^1 what my mistake.>>>

OpenStudy (anonymous):

\[ \lim_{x \rightarrow 0} sinx^{tanx}\]

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

the answer is 1.. it mean e^0

OpenStudy (anonymous):

i got 1 as well

OpenStudy (anonymous):

how it is. the question ask you to use L'Hopital..

OpenStudy (anonymous):

well hopefully somebody can explain how to do it by that method. I am not sure

zepdrix (zepdrix):

You get an indeterminate form if you plug 0 in. 0^0 I think. Hmm let's try using L'Hop. To do so we'll have to revisit an old trick.\[\huge \lim_{x \rightarrow 0} \sin x^{\tan x} \qquad = \qquad \lim_{x \rightarrow 0}e^{\ln(\sin x^{\tan x})}\]

zepdrix (zepdrix):

Understand what I did there? I exponentiated AND took the log of the problem, which didn't change it's value.

zepdrix (zepdrix):

We want to ignore the e for right now, we'll just look at the inside part. Using a familiar rule of logarithms we can write it as,\[\huge \lim_{x \rightarrow 0} \color{#CC0033}{\tan x} \color{#3366CF}{\ln(\sin x)}\]

zepdrix (zepdrix):

Confused about anything yet? D:

OpenStudy (anonymous):

ok.. then. for the left side ln(y)=tanxln(sinx)

zepdrix (zepdrix):

Sure if you want to do it that way, that works also :)

OpenStudy (anonymous):

so can i chnge tanx =SINX/cosx

zepdrix (zepdrix):

let's not. All we've done so far is rearrange things a bit.. we should STILL be getting an indeterminate form. As we approach 0 now, we'll see that our limit is aproaching \(0( -\infty)\)

zepdrix (zepdrix):

We need the indeterminate form, \(\huge \frac{0}{0}\) or \(\huge \frac{\infty}{\infty}\) in order to apply L'Hop Rule.

zepdrix (zepdrix):

So from here we can ummm take advantage of a little trick. Let's rewrite our Tangent function as \(\huge \frac{1}{cot x}\)

zepdrix (zepdrix):

\[\huge \lim_{x \rightarrow 0}\frac{\color{#3366CF}{\ln(\sin x)}}{\color{#CC0033}{\cot x}}\]

zepdrix (zepdrix):

Does that make sense how we changed it to cotangent? :o

OpenStudy (anonymous):

f'(x) = cosx/sinx g,(x)=-cosec^2(x)

zepdrix (zepdrix):

Yah looks good so far c:

zepdrix (zepdrix):

\[\large \lim_{x \rightarrow 0} \frac{\frac{\cos x}{\sin x}}{-\csc^2 x} \quad = \quad \lim_{x \rightarrow 0}\frac{-\cos x \sin^2 x}{\sin x}\] Looks like we can simplify it down a little bit.

OpenStudy (anonymous):

can x/0 because sin 0 = 0

zepdrix (zepdrix):

You need to simplify before plugging 0 in :O

zepdrix (zepdrix):

Cancel stuff out.

OpenStudy (anonymous):

ok.. so no more sin xbecause we cancel each other..

OpenStudy (anonymous):

what do you understand about the convergent and divergent???

zepdrix (zepdrix):

\[\large \lim_{x \rightarrow 0} -\cos x \cdot \sin x\]And I think this should give us a nice answer which is NOT of indeterminate form. If you're unsure, you could also apply the sine double angle to simplify it down alittle further.

zepdrix (zepdrix):

Likes with power series? I'm not very good with those to be honest :C

OpenStudy (anonymous):

as know convergent is lim> infinite is =0 if not zero it is divergent>>